# LiFT: A Surprisingly Simple Lightweight Feature Transform for Dense ViT Descriptors

### v1

@@ 2.1 Vision Transformers
Vision Transformers (ViTs) [16] have gained wide popularity as general-purpose models for multiple computer vision tasks such as image classification [16, 43, 1], object detection [54, 28, 31], segmentation [36, 40], video classification [3, 4], and more. Many of these methods adapt ViTs to different tasks by making suitable changes to the output heads [28, 51]. There have also been multiple variants of ViT like Swin [31], MViT [20], and PVT [47] which incorporate hierarchy and multiscale learning. Additionally, there are some works which try to bridge the gap between transformers and CNNs by incorporating convolutions into ViT architectures [50, 17]. In this work, we focus on improving traditional or “Plain” ViT backbones, as they are the most general and widely adopted form of ViT. These models learn powerful representations, but they suffer in terms of feature resolution, an issue which we aim to address with LiFT.
2.2 Supervision Strategies for ViTs
Many works have proposed self-supervised tasks to learn meaningful representations with ViTs. These approaches do not require labels and can make use of large amounts of unlabeled data. These self-supervised models have powerful off-the-shelf features and can act as good pre-trained initializations for finetuning on downstream tasks. Among the most popular are approaches like Momentum Contrast methods (MoCo)  [7, 24, 9] which enforce consistency between features for different augmentations of the same image. DINO [5] also utilizes a similar self-supervised approach, but uses different crops along with other augmentations. MAE [23] learns to predict masked-out image regions, and there are works like CLIP [34] which match text and image embeddings to learn semantically rich features without direct label-level supervision. Prior works [46, 21, 38] have shown that these differences in pre-training lead to significant differences in the properties of the learned features. We show that LiFT can improve the quality and usefulness of ViT features for a range of different pretraining methods.
2.3 Feature Densification
@@
Other works, like [6] and [53], have applied student/teacher distillation with feature super-resolution to improve CNN classification performance on low-resolution images. While we also perform feature super-resolution, the aim of our work is not to distill a student network, but rather we aim to generate densified features while completely avoiding finetuning the ViT backbone. [2] proposes a simpler technique to increase the density of ViT features by reducing the stride during initial image patch extraction. This does not require any training, but also becomes computationally expensive as the number of tokens increase quadratically requiring more GPU memory and FLOPs. Our LiFT approach can be thought of as a “shortcut” to achieve the benefits of denser features for a fraction of the computation cost of increased tokens.
@@ 2.4 Finetuning ViTs for Dense Tasks
The prior works most similar to LiFT are those that apply self-supervised finetuning to pretrained ViTs to improve their features for dense tasks. SelfPatch [55] and Leopart [52] both use pretrained DINO models as a starting point and improve their patch-level representations using dense self-supervised tasks. SelfPatch learns by enforcing similarity to neighboring patch features, and Leopart uses spatially dense clustering to enforce similarity within clusters to learn better part representations. These approaches finetune the full backbone with specialized training strategies and losses, which can still be expensive to train and not easily extendable to other backbones. In contrast, our approach does not require backbone finetuning at all, and instead trains a lightweight post-processing module. LiFT can be easily trained on ImageNet [14] in a self-supervised manner in very few epochs and afterwards it can generalize to multiple downstream tasks.
@@ 3.2 Lightweight Feature Transform (LiFT)
Our proposed approach, LiFT, builds on the hypothesis that, even though the ViT feature descriptors have a low spatial resolution, their high dimensionality allows them to store rich information about the image structure. This hypothesis is further supported by works on internal learning of images [39, 22, 56]. However, unlike internal learning, we propose to train a general-purpose, lightweight upsampling network using only self-supervision to double the resolution of feature descriptors obtained from a frozen, pretrained ViT. Furthermore, we can gain additional fine-level information directly from the original image at the same resolution used to generate the ViT features. We fuse these two information sources to create our final LiFT module, as illustrated in Figure 2.
Given ViT descriptors of size $\frac{H}{P}\times\frac{W}{P}$, a single LiFT expansion block scales them to $\frac{2H}{P}\times\frac{2W}{P}$ in a single forward pass. Our LiFT block is built following a U-Net-style structure [37] with skip connections, where semantically rich but coarse ViT features are combined with shallow but dense image features derived from a second input with the original image. The image input is processed through a series of convolution blocks and the resulting features are concatenated to the ViT features. Then, a single transpose convolution block is applied to generate the upscaled semantically rich features. Thanks to its fully convolutional nature, the LiFT block can be applied to any image size. One can even apply a LiFT block multiple times to further upscale the features, as shown in Section 6.4.
@@ 3.3 Training Objective
Given an image $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ and a pretrained, frozen ViT model $\mathcal{F}$ of stride $P$, we extract its features from the last layer such that $\mathcal{F}(\mathbf{x})\in\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}\times D}$, where $D$ is the feature dimension. The LiFT block $\boldsymbol{\Theta}$ upscales the features from resolution $\frac{H}{P}\times\frac{W}{P}$ to $\frac{2H}{P}\times\frac{2W}{P}$. For training, we propose the following multi-scale reconstruction objective:
@@ 3.4 Training Details
Training LiFT is fast and efficient, as it is lightweight and does not require propagating gradients through the ViT backbone. The LiFT module, $\boldsymbol{\Theta}$, is a small network with $1.1M$ trainable parameters. We train LiFT for 5 epochs on the ImageNet dataset on a single GPU. We use a learning rate of 0.001 with a batch size of 256 for stride 16 training. We use DINO [5] ViT-S/16 as our base ViT, and we apply color jitter as the only augmentation. In total, training takes only ${\sim}8$ hours on one RTX A6000 GPU. Once our LiFT module is trained, it is a general purpose feature enhancement module that can be directly applied to a range of downstream tasks without need for any further finetuning of LiFT or the frozen ViT backbone. We demonstrate the flexibility of LiFT by applying features from the same DINO+LiFT model to several tasks in Sections 4.1, 4.2, and 4.3. We present additional analysis of different design choices for both training and inference in Appendix 0.A.
@@ 3.5 Using LiFT with Downstream Modules
Our LiFT module can be easily applied to downstream tasks that directly operate on the output image features. However, LiFT can also be applied in circumstances where additional downstream modules follow the feature extracting backbone. To demonstrate this, we show how to apply LiFT to the ViTDet architecture [28], which combines a ViT backbone with a Mask-RCNN-style head [25], to perform COCO Object Detection and Segmentation. To achieve this, we inject the LiFT module after the backbone and before the Mask-RCNN head in a pretrained ViTDet model. We train this LiFT module on COCO training data using the same self-supervised objective described in Section 3.3. One can then optionally finetune the downstream head on the LiFT-generated features. We show that finetuning the head is preferable, but even without finetuning, adding LiFT is beneficial. We present these results in Section 4.4, and we provide additional details of our ViTDet+LiFT architecture in Appendix 0.B.
@@ 3.6 Baseline Methods
We propose LiFT as a self-supervised, task-agnostic, general enhancement for ViT features. For baselines, we compare with the most similar prior works, SelfPatch [52] and Leopart [55], both of which finetune the entire DINO model on custom self-supervised losses to improve the quality of the spatial features for dense tasks. We demonstrate the effectiveness of LiFT compared with these methods for a range of dense and local tasks in Section 4. In addition, we compare with the reduced token stride strategy of [2] and show the improved computational efficiency of LiFT in Section 5. Note that LiFT represents an orthogonal direction of ViT improvement that can be used in combination with methods like [2] and [55], which we show in the following sections and in Appendix 0.C.
@@ 4 Performance Benefits of LiFT
Table 1: Comparison between LiFT and other baselines on the keypoint correspondence task on SPair-71k. We report PCK@0.1 and 0.05 at multiple input resolutions.
PCK@0.1 PCK@0.05
Method/Resolution 56 112 224 448 56 112 224 448
DINO 2.04 12.67 24.76 28.6 0.51 3.61 9.54 15.33
Leopart 2.35 11.2 23.33 26.54 0.6 3.22 8.9 12.26
SelfPatch 2.13 12.18 23.03 27.34 0.44 3.61 9.32 14.44
DINO+LiFT 5.05 17.72 28.68 31.38 1.19 6.29 14.72 18.90
4.1 SPair Keypoint Correspondence
This task involves matching keypoints between pairs of images in the SPair-71k dataset. We follow the evaluation protocol of Amir et al. [2] and report Percentage of Correct Keypoints (PCK) as the metric. We extract dense features using the frozen DINO+LiFT combination trained in Section 3.4. For brevity, we report results at PCK thresholds of $0.1$ and $0.05$ in the main paper but we can also see consistent improvements at the $0.01$ threshold, as shown in Appendix 0.D. For this task, the features of all methods are bilinearly interpolated to match the original image resolution before feature matching begins. Table 1 presents the results. Compared to both the base DINO model and finetuned approaches, LiFT performs the best across all resolutions for both PCK@0.1 (left) and PCK@0.05 (right). For lower resolutions like $56\times 56$, LiFT more than doubles the performance on both metrics. Also note that under the PCK@0.1 metric, DINO+LiFT at $224\times 224$ resolution beats all non-LiFT approaches run at the higher resolution of $448\times 448$, even though both configurations produce final features of the same density.
4.2 DAVIS Video Object Segmentation
@@
Table 2: Comparison on the DAVIS video object segmentation task. We report results for the J and F mean.
Table 3: When added to ViTDet, LiFT boosts detection and segmentation performance on COCO. ${}^{\star}$ shows results without fine-tuning the Mask-RCNN head.
Method/Resolution 56 112 224 448
@@
DINO+Bilinear 10.8 23.7 37.0 53.0
DINO+LiFT 13.0 28.0 44.3 61.1
Detection Segmentation
@@
ViTDet 39.50 61.56 42.23 37.81 60.47 39.97
${}^{\star}$ 42.77 61.05 46.60 38.18 58.28 40.96
ViTDet+LiFT 45.98 66.41 49.87 40.78 63.34 43.57
Table 3: When added to ViTDet, LiFT boosts detection and segmentation performance on COCO. ${}^{\star}$ shows results without fine-tuning the Mask-RCNN head.
4.3 Unsupervised Object Discovery
We test the benefits of LiFT for Unsupervised Single Object Discovery on PASCAL VOC 2007 [18], PASCAL VOC 2012 [19], and COCO20K [30]. We apply TokenCut [48], which performs Graph Cut on features, to LiFT and the other baseline methods. Similar to prior works [48, 11, 15, 49, 35, 45] we report the Correct Localization (CorLoc) metric, which is computed as the fraction of the images in which at least one object box prediction has an IoU greater than a threshold (0.5) with a ground truth box. As shown in Table 4, LiFT gives a good boost in CorLoc, with gains across all resolutions and with LiFT outperforming the other approaches. It should be noted that the performance gains at lower resolutions are especially large. For example, at the $56\times 56$ resolution, we see an improvement of $15.8$, $17.29$, and $11.44$ on CorLoc for VOC07, VOC12, and COCO20K respectively.
Dataset Method Resolution
56 112 224 448
VOC07 DINO 20.74 50.07 65.60 68.27
Leopart 18.92 32.59 51.59 48.79
SelfPatch 18.04 41.99 62.40 63.62
DINO+LiFT 36.54 62.02 68.79 69.65
@@ 4.3 Unsupervised Object Discovery
Leopart 22.44 37.41 55.74 54.40
SelfPatch 20.19 47.32 68.02 66.48
DINO+LiFT 40.56 66.21 70.91 71.71
[…]
@@ 5 Computational Efficiency of LiFT
We have shown that feature densification with LiFT provides significant performance benefits in several tasks. We now show why LiFT is a Lightweight transform, as it is vastly more computationally efficient than other alternatives for feature densification. A trivially easy way to boost the density of ViT features is to increase the resolution of the input image, which increases token density but also computational cost. Another option is the dense feature extraction strategy of [2] which increases the number of tokens in the network by reducing the stride during patch extraction. We present a comprehensive compute cost vs. performance benefit analysis for LiFT and alternative methods, and we show that LiFT acts as a “shortcut” to achieve higher resolution features for minimal extra compute cost. We also show that LiFT can be combined with these methods for further improved performance in exchange for higher compute.
Table 5: FLOPs comparison for a single forward pass across resolutions and strides for base DINO and LiFT. Along with FLOPS, we report the performance on both SPair-71k and DAVIS.
@@
Figure 3: Performance vs. Compute Cost trade-off curve for SPair-71k keypoint correspondence. For any given FLOP-budget, DINO+LiFT achieves far superior performance. [Uncaptioned image]
5.1 LiFT, Resolution, and Stride
For this analysis, we take SPair Keypoint Correspondence and DAVIS Video Segmentation as the tasks, and we measure compute cost in FLOPs (G) against performance both with and without LiFT over a range of input resolutions and strides. As shown in rows 1 and 2 of Table 5, LiFT introduces a small increase in FLOPs from $4.34$G to $5.30$G ($+22\%$) but also brings a significant performance improvement of $3.92$, $5.18$, and $11.3$ points on the PCK@0.1, PCK@0.05, and J and F Mean metrics respectively. For comparison, reducing the stride from $16$ to $8$ (row 1 vs. row 3) gives a similar level of improvement, but nearly quadruples the FLOPs from $4.34$G to $16.07$G ($+270\%$). A similar trend of large improvements can be seen when comparing pairs of rows (3 vs. 4, 5 vs. 6) with DINO vs. DINO+LiFT. The best overall results are achieved by combining LiFT with increased resolution and reduced stride, but this also comes with the highest computation cost. It should also be noted that DINO+LiFT at $224\times 224$ resolution and stride 8 (row 4) uses 3×3$\times$3 × less FLOPs than DINO at $448\times 448$ resolution at stride 8 (row 7) while having similar performance on the PCK metrics.
5.2 Cost vs. Performance Trade-off Curve
Next, we present a comprehensive Compute Cost vs. Performance trade-off analysis on SPair Keypoint Correspondence. By incrementally increasing the input resolution, we can gradually increase both the performance and inference cost for DINO+LiFT and the baseline methods DINO, Leopart, and Selfpatch. As shown in Figure 5, we find that LiFT significantly outperforms all other methods at any given FLOP allowance, in most cases seeing a ${\sim}20\%$ performance gain. Alternatively, LiFT can be run at a lower input resolution to achieve equivalent performance at a fraction of the compute cost. For example, to achieve a score of $15.0$ in PCK@0.05, DINO+LiFT only requires ${\sim}6.25$ Giga-FLOPs of compute power, while the other baselines require at least $15$ Giga-FLOPs. At any point on the trade-off curve, DINO+LiFT far surpasses the alternatives.
@@ 5.1 LiFT, Resolution, and Stride
5.3 Parameter Count
We acknowlege that the addition of the LiFT module slightly increases the overall model size and parameter count as shown in Table 7. However, this addition is quite small and only represents a $+5.7\%$ change in total parameters. For comparison, the jump from ViT-S to ViT-B results in a $+304\%$ increase in parameters. Furthermore, for dense tasks like SPair Keypoint Correspondence, we find that the performance benefits provided by LiFT far exceed the benefits of a larger backbone. Note that the methods Leopart and SelfPatch do not increase the parameter count of the ViT, as they finetune the DINO backbone instead of introducing new modules. However, we believe the major performance benefits of LiFT strongly justify the small extra costs in Parameter Count and Inference FLOPs for a given resolution. And as shown in the previous section, for any fixed FLOP-budget, DINO+LiFT achieves far superior performance.

### v2

@@ 2.1 Vision Transformers
Recently, Vision Transformers (ViTs) [18] have gained wide popularity as general-purpose models for multiple computer vision tasks such as image classification [18, 51, 1], object detection [63, 33, 36], segmentation [44, 48], video classification [3, 4], and more. Many of these methods adapt ViTs to different tasks by making suitable changes to the output heads [33, 60]. There have also been multiple variants of ViT like Swin [36], MViT [22], and PVT [56] which incorporate hierarchy and multiscale learning. Additionally, there are some works which try to bridge the gap between transformers and CNNs by incorporating convolutions into ViT architectures [59, 19]. In this work, we focus on improving traditional or “Plain” ViT backbones, as they are the most general and widely adopted form of ViT. These models learn powerful representations, but they suffer in terms of feature resolution, an issue which we aim to address with LiFT. Many self-supervised ViT pretraining methods have been proposed to learn powerful off-the-shelf features, such as DINO [6], MoCo [8, 28, 10], and MAE [27]. Prior works [54, 25, 46] have shown that different pretraining leads to significant differences in the properties of the learned features. We show that LiFT can improve the quality and usefulness of ViT features for a range of pretraining methods.
2.2 Extracting Denser Features from ViTs
@@
2.3 Feature Upsampling
Over the years, many works have been proposed to convert lower resolution features into higher resolution ones. This includes basic methods like bilinear interpolation, and Joint Bilinear Upsampling (JBU) [31], which leverages high-frequency information from the input image to improve bilinear interpolation. Our LiFT module leverages the input image similarly. Resize-convolutions [41] are proposed as an alternative to deconvolution. Many more sophisticated upsampling operators have been proposed [55, 39, 38, 37, 15], however these methods are designed to be trained in the context of their particular architecture and downstream task, unlike our LiFT module, which is stand-alone and task-agnostic. Moreover, many of these methods are designed to be integrated with convolutional encoder/decoder architectures, and thus are not easily applied to ViT backbones. [50] proposes a GAN-based approach for CNN feature densification which requires careful training and a mixture of adversarial and focal loss. In comparison, our LiFT approach is easy to train through a simple self-supervised objective. Other works, like [7] and [62], have applied student/teacher distillation with feature super-resolution to improve CNN classification performance on low-resolution images. While we also perform feature super-resolution, the aim of our work is not to distill a student network, but rather we aim to generate densified features while avoiding finetuning the ViT backbone.
Finally, we would like to acknowledge a concurrently published work, FeatUp [23], which presents two feature upsampling methods, one based on JBU and a second based on implicit network learning. Like LiFT, FeatUp trains general-purpose stand-alone feature upsamplers for multiple backbones and downstream tasks, though its learning process is more complex, as it requires training an additional downsampler module which is later discarded. LiFT and FeatUp${}_{\text{(JBU)}}$ are both feedforward, making them easy to apply to any image, though we believe LiFT’s feature maps are sharper and better matched to the object boundaries than the JBU-enhanced features. FeatUp${}_{\text{(Implicit)}}$ produces sharp, impressive feature maps, though compared with LiFT and FeatUp${}_{\text{(JBU)}}$, it is a much less scalable method, as it trains a new implicit network for every input image. Overall, we believe LiFT captures the best properties of both FeatUp variations.
@@ 2.4 Finetuning ViTs for Dense Tasks
Another line of research has focused on finetuning self-supervised ViT backbones to make their features better suited to dense, locally-focused tasks. SelfPatch [64] and Leopart [61] both use pretrained DINO models as a starting point and improve their patch-level representations using dense self-supervised tasks. SelfPatch learns by enforcing similarity to neighboring patch features, and Leopart uses spatially dense clustering to enforce similarity within clusters to learn better part representations. These approaches finetune the full backbone with specialized training strategies and losses, which can still be expensive to train and not easily extendable to other backbones. In contrast, our approach does not require backbone finetuning at all, and instead trains a lightweight post-processing module. LiFT can be easily trained on ImageNet [16] in a self-supervised manner in very few epochs and afterwards it can generalize to multiple downstream tasks.
@@ 3.2 Lightweight Feature Transform (LiFT)
Our proposed approach, LiFT, builds on the hypothesis that, even though the ViT feature descriptors have a low spatial resolution, their high dimensionality allows them to store rich information about the image structure. This hypothesis is further supported by works on internal learning of images [47, 26, 65]. However, unlike internal learning, we propose to train a general-purpose, lightweight upsampling network using only self-supervision to double the resolution of feature descriptors obtained from a frozen, pretrained ViT. Furthermore, we can gain additional fine-level information directly from the original image at the same resolution used to generate the ViT features. We fuse these two information sources to create our final LiFT module, as illustrated in Figure 2. Given ViT descriptors of size $\frac{H}{P}\times\frac{W}{P}$, a single LiFT expansion block scales them to $\frac{2H}{P}\times\frac{2W}{P}$ in a single forward pass. Our LiFT block is built following a U-Net-style structure [45] with skip connections, where semantically rich but coarse ViT features are combined with shallow but dense image features derived from a second input with the original image. The image input is processed through a series of convolution blocks and the resulting features are concatenated to the ViT features. Then, a single transpose convolution block is applied to generate the upscaled semantically rich features. Thanks to its fully convolutional nature, the LiFT block can be applied to any image size. One can even apply a LiFT block multiple times to further upscale the features, as shown in Section 6.
@@ 3.3 Training Objective
Given an image $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ and a pretrained, frozen ViT model $\mathcal{F}$ of stride $P$, we extract last-layer features $\mathcal{F}(\mathbf{x})\in\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}\times D}$, where $D$ is the feature dimension. The LiFT block $\boldsymbol{\Theta}$ upscales the features from $\frac{H}{P}\times\frac{W}{P}$ to $\frac{2H}{P}\times\frac{2W}{P}$. For training, we propose the following multi-scale reconstruction objective:
@@ 3.4 Training Details
Training LiFT is fast and efficient, as it is lightweight and does not require propagating gradients through the ViT backbone. The LiFT module, $\boldsymbol{\Theta}$, is a small network with $1.2M$ trainable parameters. We train LiFT for 5 epochs on the ImageNet dataset on a single GPU. We use a learning rate of 0.001 with a batch size of 256 for stride 16 training. We use DINO [6] ViT-S/16 as our base ViT, and we apply color jitter as the only augmentation. In total, training takes only ${\sim}8$ hours on one RTX A6000 GPU. Once our LiFT module is trained, it is a general purpose feature enhancement module that can be directly applied to a range of downstream tasks without need for any further finetuning. We demonstrate the flexibility of LiFT by applying features from the same DINO+LiFT model to several tasks in Section 4. We present additional analysis of different design choices for training and inference in Appendix A.
@@ 3.5 Using LiFT with Downstream Modules
Our LiFT module can be easily applied to downstream tasks that directly operate on the output image features. However, LiFT can also be applied in circumstances where additional downstream modules follow the feature extracting backbone. To demonstrate this, we show how to apply LiFT to the ViTDet architecture [33], which combines a ViT backbone with a detector head to perform COCO Object Detection and Segmentation. To achieve this, we inject the LiFT module after the backbone and before the head in a pretrained ViTDet model. We train this LiFT module on COCO training data using the same self-supervised objective described in Section 3.3. We then briefly finetune the pretrained head on the LiFT features. We present these results in Section 4, and we provide additional details of our ViTDet+LiFT architecture in Appendix B.
@@ 3.6 Baseline Methods
In addition to the base DINO features, we compare to three types of baselines. The first are baselines which finetune DINO features to enhance them without increasing their density. This includes Leopart [64] and SelfPatch [61]. The second are basic feature upsampling methods, which includes bilinear interpolation (BL), resize-convolutions (RC), and Joint Bilinear Upsampling (JBU). For these methods, we double the feature resolution to match LiFT. We use the generalized JBU implementation of [23], which includes learnable parameters. We train both JBU and RC following the same optimization protocol as our LiFT module. The final type are dense ViT feature extraction methods, specifically increased image resolution and reduced stride [2]. Note that LiFT represents an orthogonal direction of ViT improvement that can be used in combination with methods like [2] and [64], which we show in the following sections and in Appendix C.
@@ 4 Performance Benefits of LiFT
Table 1: Comparison between LiFT and other baselines on the Keypoint Correspondence task on SPair-71k. We report PCK@0.1/0.05/0.01 at multiple input resolutions.
PCK@0.1 PCK@0.05 PCK@0.01
Method/Res. 56 112 224 448 56 112 224 448 56 112 224 448
DINO 2.04 12.67 24.76 28.6 0.51 3.61 9.54 15.33 0.01 0.2 0.54 1.4
Leopart 2.35 11.2 23.33 26.54 0.6 3.22 8.9 12.26 0.05 0.1 0.47 0.79
@@
DINO+RC 3.87 14.02 26.09 30.21 0.94 4.41 11.51 16.76 0.03 0.16 0.51 1.7
DINO+JBU 3.08 12.26 24.87 29.11 0.82 3.82 10.6 16.01 0.03 0.17 0.47 1.42
[…]
SPair Keypoint Correspondence. This task involves matching keypoints between pairs of images in the SPair-71k dataset. We follow the evaluation protocol of Amir et al. [2] and report Percentage of Correct Keypoints (PCK) with three different distance thresholds. We extract dense features using the frozen DINO+LiFT combination trained in Section 3.4. For this task, the features of all methods are resized to match the original image resolution using Lanczos interpolation before feature matching begins. Table 1 presents the results. Compared to both the base DINO model and other baselines, LiFT performs the best across all resolutions for all metrics. For lower resolutions like $56\times 56$, LiFT more than doubles the performance of DINO. Also note that under the PCK@0.1 metric, DINO+LiFT at $224\times 224$ resolution beats DINO run at $448\times 448$, even though both configurations produce final features of the same density.
[…]
Table 2: DAVIS Video Object Segmentation results for the J & F Mean metric.
Method/Res. 56 112 224 448
@@
DINO+RC 11.0 24.0 37.4 53.2
DINO+JBU 11.2 26.2 39.0 54.5
@@
Table 3: COCO20K Unsupervised Object Discovery results for CorLoc metric.
Method/Res. 56 112 224 448
DINO 16.28 40.08 53.98 57.99
@@
DINO+RC 22.92 42.53 54.52 58.40
DINO+JBU 21.36 43.87 55.45 58.82
[…]
Table 4: COCO Detection and Segmentation results with ViTDet+LiFT for two different heads: Mask R-CNN (MR) and Cascade R-CNN (CR).
Detection Segmentation
Method AP AP50 AP75 AP AP50 AP75
ViTDet(MR) 39.50 61.56 42.23 37.81 60.47 39.97
ViTDet(MR)+LiFT 45.98 66.41 49.87 40.78 63.34 43.57
[…]
ViTDet(CR)+LiFT 47.92 65.88 51.67 41.42 63.21 44.45
Unsupervised Object Discovery. We test the benefits of LiFT for Unsupervised Single Object Discovery on COCO20K [35]. We apply TokenCut [57], which performs Graph Cut on features, to LiFT and the other baseline methods. Similar to prior works [57, 12, 17, 58, 43, 53] we report the Correct Localization (CorLoc) metric, which is computed as the fraction of the images in which at least one object box prediction has an IoU greater than a threshold (0.5) with a ground truth box. As shown in Table 3, LiFT gives a good boost in CorLoc, with gains across all resolutions and with LiFT outperforming the other approaches. It should be noted that the performance gains at lower resolutions are especially large. For example, at the $56\times 56$ resolution, we see an improvement of $11.44$ over DINO. In Appendix D, we present additional results on PASCAL VOC 2007 [20] and PASCAL VOC 2012 [21], and we observe similar overall trends.
COCO Detection and Segmentation. To further demonstrate the versatility of LiFT, we show how it can be applied to COCO Detection and Segmentation using a ViTDet+LiFT model as described in Section 3.5. We present results for two ViTDet variants using different heads: Mask R-CNN (MR) [29], and Cascade R-CNN (CR) [5]. In Table 4, ViTDet(MR)+LiFT gives a ${\sim}6.5\%$ boost for Detection AP and a ${\sim}3\%$ boost for Segmentation AP. Smaller but still consistent boosts are also achieved when combining LiFT with ViTDet(CR).
@@ 5 Computational Efficiency of LiFT
We have shown that feature densification with LiFT provides significant performance benefits in several tasks. We now show why LiFT is a Lightweight transform, as it is vastly more computationally efficient than other alternatives that increase the token density of the ViT backbone. A trivially easy way to boost the density of ViT features is to increase the resolution of the input image. Another option is the dense feature extraction strategy of [2] which increases the number of tokens in the network by reducing the stride during patch extraction. Both of these methods increase token density but at a high computational cost. We present a comprehensive compute cost vs. performance benefit analysis for LiFT against and in combination with these methods, and we show that LiFT acts as a “shortcut” to achieve higher resolution features for minimal extra compute.
Table 5: FLOPs comparison for a single forward pass across resolutions and strides for DINO and DINO+LiFT. Along with FLOPS, we report the performance on both SPair-71k and DAVIS.
@@
Figure 3: Performance vs. Compute Cost trade-off curve for SPair-71k keypoint correspondence. For any given FLOP-budget, DINO+LiFT achieves far superior performance.
LiFT, Resolution, and Stride. For this analysis, we take SPair Keypoint Correspondence and DAVIS Video Segmentation as the tasks, and we measure compute cost in FLOPs (G) against performance both with and without LiFT over a range of input resolutions and strides. As shown in rows 1 and 2 of Table 5, LiFT introduces a small increase in FLOPs from $4.34$G to $5.30$G ($+22\%$) but also brings a significant performance improvement of $3.92$ and $11.3$ points on the PCK@0.1 and J & F Mean metrics respectively. For comparison, reducing the stride from $16$ to $8$ (row 1 vs. row 3) gives a similar level of improvement, but nearly quadruples the FLOPs from $4.34$G to $16.07$G ($+270\%$). A similar trend of large improvements can be seen when comparing pairs of rows (3 vs. 4, 5 vs. 6) with DINO vs. DINO+LiFT. The best overall results are achieved by combining LiFT with increased resolution and reduced stride, but this also comes with the highest computation cost. It should also be noted that DINO+LiFT at $224\times 224$ resolution and stride 8 (row 4) uses 3×3$\times$3 × less FLOPs than DINO at $448\times 448$ resolution at stride 8 (row 7) while having similar PCK performance.
Performance Trade-Off Curve. We present a comprehensive Compute Cost vs. Performance trade-off curve for SPair Keypoint Correspondence in Figure 5. By incrementally increasing the input resolution, we can gradually increase both the performance and inference cost for DINO and DINO+LiFT features. We find that LiFT significantly outperforms the base DINO features at any given FLOP allowance, in most cases seeing a ${\sim}20\%$ performance gain. Alternatively, LiFT can be run at a lower input resolution to achieve equivalent performance at a fraction of the compute cost. For example, to achieve a score of $15.0$ in PCK@0.05, DINO+LiFT only requires ${\sim}6.25$ Giga-FLOPs of compute power, while the standard DINO requires at least $15$ Giga-FLOPs.
@@
Parameter Count. We acknowlege that the addition of the LiFT module slightly increases the overall model size and parameter count as shown in Table 7. However, this addition is quite small and only represents a $+5.7\%$ change in total parameters. For comparison, the jump from ViT-S to ViT-B results in a $+304\%$ increase in parameters. Furthermore, for dense tasks like SPair Keypoint Correspondence, we find that the performance benefits provided by LiFT far exceed the benefits of a larger backbone. Note that the methods Leopart and SelfPatch do not increase the parameter count of the ViT, as they finetune the DINO backbone instead of introducing new modules. We believe the major performance benefits of LiFT justify the small extra costs in Parameter Count and Inference FLOPs for a given resolution. And, as shown in the previous section, for any fixed FLOP-budget, DINO+LiFT achieves far superior performance.